If the ratio of the width of a rectangle to its length is 7/8, and the length is 40 centimeters, find the width
Answer by jim_thompson5910(35256)

W/L = 7/8

W/40 = 7/8

W = 40*(7/8)

W = 35

The width is 35 cm

Let me know if you need more help or if you need me to explain a step in more detail.
Feel free to email me at jim_thompson5910@hotmail.com
or you can visit my website here: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim